# Physics-Constrained Learning of Energy-Preserving Stencils for Maxwell’s Equations

### v3

@@ Abstract
We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system. Using high-fidelity training data from a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator in Maxwell’s equations. We formulate a convex quadratic program for the stencil coefficients with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol. We compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned operators in time-dependent Maxwell simulations using a Crank–Nicolson (CN) discretization. Numerical experiments, including cases with nonstandard target operators and noisy training data, show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point solvers produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
@@ 1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [21]. Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions. Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, for example via regression or neural networks. Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
Most existing data-driven schemes aim to minimize approximation error but do not explicitly enforce structural properties such as stability, conservation, or energy balance. For Maxwell’s equations this is a significant limitation, since standard FDTD schemes are carefully designed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level.
@@
represented as a convolution stencil, directly from data. We impose linear constraints on the stencil that enforce skew-adjointness of $D$ with respect to the discrete $L^{2}$ inner product. This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, regardless of the details of the training data.
Classical FDTD schemes such as the Yee method [21] are hand-derived so that the discrete curl operator is skew-adjoint, yielding exact semi-discrete energy conservation. High-order extensions and operator-splitting schemes similarly maintain this structure [3, 6].
Data-driven PDE schemes learn local stencils or coarse-grid operators directly from numerical data. Bar-Sinai et al. learn finite-difference discretizations for Burgers and Navier–Stokes, showing that learned operators can match or exceed classical schemes on specific training distributions [1]. Many follow-up works learn PDE operators or closure models [17, 11, 20, 18, 9], but typically do not enforce stability, skew-adjointness, or conservation laws.
@@
Recent neural-operator surrogates for electromagnetics incorporate physics through architectural biases or penalty terms. Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material dispersion structure [14]. Domain-decomposition neural operators couple learned subdomain solvers through interface penalties or consistency conditions [15, 8]. Other neural operator frameworks embed conservation via soft constraints or corrective layers [13, 12]. These approaches improve stability but enforce physical structure only approximately, through optimization.
@@
Building on this context, our main contributions are as follows. First, we formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data, with an objective given by a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, and linear constraints that encode skew-symmetry of the stencil coefficients together with simple box bounds. Second, we show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity, yielding what is, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law. Third, we characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$. Finally, we implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations; the experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
Although our experiments focus on a simple 1D periodic model, the framework and energy constraints extend naturally to higher-dimensional settings and more complex material models. Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation.
Our method imposes hard linear constraints that make the learned convolution stencil exactly skew-adjoint, guaranteeing a discrete electromagnetic energy identity independent of training accuracy or solver choice. This places our work closest to Bar-Sinai et al. but with strict structure preservation rather than penalty-based regularization, and complementary to physics-informed neural operators, which typically enforce conservation approximately. To the best of our knowledge, this is the first work to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol. This provides a bridge between unconstrained data-driven discretizations and classical energy-preserving FDTD schemes, while remaining compatible with standard convex optimization toolchains (projected gradient, Nesterov acceleration, ADMM, and interior-point solvers).
Beyond the homogeneous model, Sections 6.6 and 6.7 illustrate how the same constrained learning framework can recover nonstandard, effective operators and regularize noisy or mildly inconsistent training data.
@@ 2.1 Skew-symmetric convolution operators and stencil constraints
where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$.[3]
@@
[3]Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$. Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
@@ 3 Wave speed and CFL for the learned operator
which expresses the admissible time step solely in terms of the learned symbol (12). In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.In this work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator. A systematic comparison with leapfrog using the derived CFL bound is left for future work
@@ 4.1 Training data from spectral 1D Maxwell
stores the corresponding time derivatives, with $(dU/dt)[s,0,i]=\partial_{t}E^{(s)}(x_{i})$ and $(dU/dt)[s,1,i]=\partial_{t}H^{(s)}(x_{i})$.
Here $n_{\mathrm{sims}}$ denotes the number of independent training samples (or snapshots), each consisting of a randomly generated Maxwell state $(E,H)$ and its corresponding time derivative. The spatial domain $[0,L]$ is discretized using $N$ uniformly spaced grid points $x_{i}=i\Delta x$, $i=0,\dots,N-1$, where $\Delta x=L/N$.
@@ 5.2 ADMM
To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [4], [5] scheme. We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
@@
5.3 Interior-point reference (CVXPY)
@@ 6 Numerical experiments
All experiments were implemented in Python using NumPy for linear algebra and SciPy for FFTs. The ADMM and projected-gradient/Nesterov solvers are written explicitly as described in Section 5. The interior-point reference solution is computed using CVXPY (version X.Y) with the SCS solver. Runtimes are reported for a single-threaded run on a (Apple M4, 16 GB RAM) using Python 3.11. Although absolute times depend on hardware and implementation details, the relative trends between first-order methods (PG/NAG) and the second-order ADMM/CVX approach remained consistent across several runs.
@@ 6.1 Experimental setup
$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1),$ (22)
@@ 6.2 Learned stencils for $R=1$
Table 1: Learned stencils for radius $R=1$.
In In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution for R =1. The final-time and space-time profiles in Figure 2 are visually indistinguishable from the reference solution.
Figure 1: Space-Time Evolution
@@
6.3 Error and energy residual versus stencil radius
To explore the effect of stencil size, we repeat the learning and time-integration procedure for radii $R=1,2,3,4$. Figure 3 shows the final-time relative $L^{2}$ error $\mathrm{err}(w)$ for each method as a function of $R$. We initialize all optimizations at the zero stencil; even from this uninformative start, the energy-constrained learner recovers stencils close to classical centered differences. Figure 3 shows that for $R=1$ all four solvers achieve comparable accuracy, with errors on the order of $5\times 10^{-3}$. As the radius increases, the behavior diverges: the projected-gradient and Nesterov methods exhibit substantially larger errors, with PG reaching $\mathcal{O}(1)$ error and NAG reaching $\mathcal{O}(10^{-2})$ for $R=3$. In contrast, the ADMM and CVX solutions maintain errors in a narrow band between approximately $3\times 10^{-3}$ and $6\times 10^{-3}$ for all radii tested. This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.We believe the gap for larger R is primarily due to conditioning and the fact that PG/NAG only project onto the equality constraint, whereas ADMM/CVX solve the full KKT system with box constraints. Intersection projections (e.g., Dykstra) could close this gap but would blur the simplicity advantage of PG/NAG; we leave such variants to future work
Figure 3: Final-time relative $L^{2}$ error in the electric field versus stencil radius $R$ for each solver, using Crank–Nicolson time integration.
@@ 6.4 Discrete energy conservation for learned Maxwell stencils
The phase plots in Fig. 5(b) provide a more stringent test of accuracy. For $R=1,2,3$, the dispersion curves $\arg\mu(\theta)/\theta$ of the learned operators closely track those of the corresponding centered differences over the range of resolved wavenumbers, with only small deviations at the highest modes; in some mid-frequency bands the learned stencils are slightly closer to the continuous reference. Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties. see, e.g., [21, 2]
@@ 6.6 Learning a nonstandard skew-adjoint operator
(a) Noisy Discrete Energy Simulation
(b) Noisy E(x,t)
Figure 6: Nonstandard vs Standard Skew-adjoint operator
@@ 6.7 Noisy training and ill-posed least squares
To emulate noisy or partially inconsistent experimental data, we add Gaussian noise to the time derivatives $dU/dt$ used in the regression problem (24). Specifically, we draw $\tilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d. $N(0,\sigma^{2})$ and form the corresponding least-squares system $(A,b_{\mathrm{noisy}})$. We then compute three stencils: (i) the standard centered difference $w_{\mathrm{CD}}$, (ii) an unconstrained least-squares stencil $w_{\mathrm{LS}}$ obtained by minimizing (24) with no constraints, and (iii) an energy-constrained stencil $w_{\mathrm{QP}}$ obtained by solving the quadratic program (20) with the same noisy data.
@@
Figure 7(b) shows the final-time electric field $E(x,T)$ for all three operators. The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the Central difference baseline, while the unconstrained LS stencil produces a completely corrupted field. This experiment illustrates a regime where training data are noisy or mildly inconsistent, and the energy constraints act as a form of physics-based regularization: they prevent overfitting to noise and guarantee a discrete Poynting theorem irrespective of the data quality.
(a) Noisy Discrete Energy Simulation
(b) Noisy E(x,t)
Figure 7: Noisy training and ill-posed least squares
Figure 8: Space-Time Plot of the Electric Field Under Noise Perturbation

### v4

@@ Abstract
We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system. Starting from high-fidelity spectral training data, we learn a linear convolution stencil that approximates the spatial derivative in Maxwell’s equations. The stencil coefficients are obtained from a convex quadratic program with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol. We show that the learned operator has a purely imaginary symbol, induces purely oscillatory modal dynamics, and preserves the discrete Maxwell energy exactly under Crank–Nicolson time-stepping. Numerical experiments with smooth traveling-wave solutions demonstrate that the energy-constrained, ADMM-learned stencil with radius $R=1$ achieves approximately second-order spatial convergence in the PDE error, matching the expected order of the standard centered finite-difference stencil. We compare projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point solver for the resulting quadratic program, and we illustrate extensions to nonstandard target operators and noisy training data. Overall, the learned stencils achieve accuracy comparable to classical central differences while exactly preserving a discrete electromagnetic energy, making them viable drop-in, structure-preserving spatial operators for finite-difference time-domain methods and data-driven PDE solvers.
@@ 1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [25, 26]. Classical schemes, such as the Yee method [28], are derived directly from the continuous equations: spatial derivatives are approximated using fixed central-difference stencils, and stability is enforced via analytical CFL conditions. These schemes are carefully designed so that the discrete curl operator is skew-adjoint, which yields exact semi-discrete energy conservation and compatible discrete divergence constraints.
Recently, there has been growing interest in data-driven discretizations, in which discrete operators are learned from numerical or experimental data, for example via regression or neural networks. Examples include data-driven finite-difference discretizations for Burgers’ equation and the Navier–Stokes equations [1], data-driven closures for turbulent flows [21, 14], and sparse-regression or neural approaches for discovering governing equations and hidden dynamics [24, 22, 7]. These methods typically aim to minimize approximation error on a chosen training distribution, but they do not, in general, explicitly enforce structural properties such as stability, conservation, or energy balance.
For Maxwell’s equations, this is a significant limitation. Classical FDTD schemes are constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level, ensuring that numerical solutions respect key physical invariants. Learned operators that ignore these structures risk exhibiting uncontrolled numerical dissipation or growth, even if they fit the training data well.
@@
represented as a convolution stencil, directly from data. We impose linear constraints on the stencil that enforce skew-adjointness of $D$ with respect to the discrete $L^{2}$ inner product. This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, independently of the details of the training data or the optimization procedure.
Our approach is inspired by two lines of work. On the one hand, classical structure-preserving discretizations for Maxwell and related PDEs use staggered grids and skew-adjoint discrete curl operators to obtain discrete energy identities and long-time stability [4, 8, 19]. High-order and operator-splitting variants maintain these properties by design [4, 3]. On the other hand, data-driven PDE schemes learn local stencils or coarse-grid operators from high-fidelity simulations, often achieving accuracy comparable to or better than hand-crafted schemes on specific tasks [1, 11, 6]. However, these learned operators usually encode physical structure only implicitly, through the choice of loss function, training data, or architecture, rather than via hard constraints.
@@
Related ideas also appear in recent neural-operator surrogates for electromagnetics and other time-dependent PDEs. Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material-dispersion structure [17, 13]. Domain-decomposition neural operators couple learned subdomain solvers through interface penalties or consistency conditions [18, 10]. Other frameworks embed conservation laws via soft constraints or corrective layers [16, 15, 27]. These approaches can significantly improve stability and physical fidelity, but they generally enforce conservation or energy balance only approximately, depending on optimization and hyperparameters.
@@
Building on this context, our main contributions are as follows. We formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data. The objective is a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, subject to linear constraints that encode skew-adjointness of the stencil coefficients together with simple box bounds. We show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, which in turn yields an exact semi-discrete electromagnetic energy identity and expresses the CFL condition directly in terms of the stencil’s Fourier symbol $\mu(\theta)$. This yields, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law. We characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator via this symbol. We implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations. We also illustrate how the same constrained learning framework can recover nonstandard, coarse-grid effective operators and regularize synthetic spectral training data contaminated by noise or mild inconsistencies. In particular, the ADMM-learned radius-$R=1$ stencil exhibits approximately second-order spatial convergence for smooth traveling-wave solutions, matching the expected order of the standard centered finite-difference stencil it approximates. The experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and the interior-point solver produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime. Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation. To the best of our knowledge, this is the first data-driven convolution stencil for the one-dimensional Maxwell system whose learned coefficients are constrained to guarantee a semi-discrete electromagnetic energy identity.
The paper is organized as follows. Sections 2–3 introduce the semi-discrete 1D Maxwell system, derive the corresponding discrete energy identity, and analyze the numerical wave speed and CFL restriction associated with the learned operator. Sections 4–6 describe the learning formulation and optimization algorithms, and present numerical experiments that evaluate accuracy, energy behavior, dispersion, and robustness to nonstandard targets and noisy training data.
@@ 2.1 Skew-symmetric convolution operators and stencil constraints
where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$; see, e.g., [4]. ∎
@@
Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$. Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
@@ 3 Wave speed and CFL for the learned operator
which expresses the admissible time step solely in terms of the learned symbol (12). In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator. In the present work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator. A systematic comparison with leapfrog using the derived CFL bound is left for future work.
@@ 4.1 Training data from spectral 1D Maxwell
stores the corresponding time derivatives, with $(dU/dt)[s,0,i]=\partial_{t}E^{(s)}(x_{i})$ and $(dU/dt)[s,1,i]=\partial_{t}H^{(s)}(x_{i})$. Here $n_{\mathrm{sims}}$ denotes the number of independent training samples (or snapshots), each consisting of a randomly generated Maxwell state $(E,H)$ and its corresponding time derivative. The spatial domain $[0,L]$ is discretized using $N$ uniformly spaced grid points $x_{i}=i\Delta x$, $i=0,\dots,N-1$, where $\Delta x=L/N$.
@@ 5.2 ADMM
To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [5] scheme. We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
@@
5.3 Interior-point reference
@@ 6 Numerical experiments
All experiments were implemented in Python using NumPy for linear algebra and SciPy for FFTs. The ADMM and projected-gradient/Nesterov solvers are written explicitly as described in Section 5. The interior-point reference solution is computed using CVXPY with the SCS solver. Runtimes are reported for a single-threaded run on an Apple M4 laptop with 16 GB RAM using Python 3.11. Although absolute times depend on hardware and implementation details, the relative trends between first-order methods (PG/NAG) and the second-order ADMM/CVX approach remained consistent across several runs.
@@ 6.1 Experimental setup
$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1).$ (22)
@@ 6.2 Learned stencils
Table 1: Learned stencils for radius $R=1$ compared with the exact central-difference stencil.
In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution. The final-time and space-time profiles in Figure 2 are visually indistinguishable from the reference solution.
Figure 1: Space-time evolution of the electric field $E(x,t)$ for the learned stencils and the exact central-difference stencil.
@@
6.3 Spatial convergence for the ADMM–learned stencil
We consider the 1D Maxwell system in normalized units,
$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E,$ (24)
with $\varepsilon_{0}=\mu_{0}=1$ so that the wave speed is $c=1$. On the periodic domain $x\in[0,L]$ we choose the traveling-wave exact solution
$E_{\mathrm{ex}}(x,t)=\sin\!\left(\frac{2\pi}{L}(x-t)\right),\qquad H_{\mathrm{ex}}(x,t)=\cos\!\left(\frac{2\pi}{L}(x-t)\right),$ (25)
which satisfies (24) and is $L$-periodic in space. The initial data at $t=0$ are taken from (25), and the numerical solution is advanced to a final time $T=10$ using the Crank–Nicolson time integrator with a spatial derivative given by the ADMM–learned stencil of radius $R=1$.
Let $E^{N}_{i}$ denote the numerical approximation to $E_{\mathrm{ex}}(x_{i},T)$ at the grid points $x_{i}=i\,\Delta x$, $i=0,\dots,N_{x}-1$. We measure the relative discrete $L^{2}$ error in the electric field by
$\mathrm{err}_{E}(\Delta x)=\frac{\left\|E^{N}-E_{\mathrm{ex}}(\cdot,T)\right\|_{\ell^{2}}}{\left\|E_{\mathrm{ex}}(\cdot,T)\right\|_{\ell^{2}}},\qquad\|v\|_{\ell^{2}}^{2}=\Delta x\sum_{i=0}^{N_{x}-1}v_{i}^{2}.$ (26)
The timestep is chosen proportionally to the mesh width, $\Delta t=0.2\,\Delta x$, so that we probe the spatial convergence of the learned stencil.
Table 2: Convergence of the ADMM–learned stencil with radius $R=1$. Relative $L^{2}$ error in $E(\cdot,T)$
$N_{x}$ $\Delta x$ rel. $L^{2}(E)$ error order
64 $1.562\times 10^{-2}$ $2.097\times 10^{0}$
128 $7.812\times 10^{-3}$ $4.875\times 10^{-1}$ $2.10$
256 $3.906\times 10^{-3}$ $1.192\times 10^{-1}$ $2.03$
512 $1.953\times 10^{-3}$ $585\times 10^{-2} 2.2$02.20
The table2 confirms approximately second-order spatial accuracy for the ADMM–learned radius-$1$ stencil.
Figure 3: Log–log plot of the relative $L^{2}$ error $\mathrm{err}_{E}(\Delta x)$ in $E(\cdot,T)$ versus $\Delta x$ for the ADMM–learned stencil.
@@ 6.5 Dispersion analysis
The phase plots in Fig. 5(b) provide a more stringent test of accuracy. For $R=1,2,3$, the dispersion curves $\arg\mu(\theta)/\theta$ of the learned operators closely track those of the corresponding centered differences over the range of resolved wavenumbers, with only small deviations at the highest modes; in some mid-frequency bands the learned stencils are slightly closer to the continuous reference. Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties; see, e.g., [25, 2].
@@ 6.6 Learning a nonstandard skew-adjoint operator
(a) Final-time electric field $E(x,T)$ for nonstandard and learned stencils.
(b) Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$.
Figure 6: Nonstandard vs. standard skew-adjoint operators: learned stencil $w_{\mathrm{QP}}$ compared with $w^{\star}$ and the centered difference $w_{\mathrm{CD}}$.
@@ 6.7 Noisy training and ill-posed least squares
To emulate noisy or partially inconsistent experimental data, we add Gaussian noise to the time derivatives $dU/dt$ used in the regression problem (17). Specifically, we draw $\widetilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d. $N(0,\sigma^{2})$ and form the corresponding least-squares system $(A,b_{\mathrm{noisy}})$. We then compute three stencils: (i) the standard centered difference $w_{\mathrm{CD}}$, (ii) an unconstrained least-squares stencil $w_{\mathrm{LS}}$ obtained by minimizing (17) with no constraints, and (iii) an energy-constrained stencil $w_{\mathrm{QP}}$ obtained by solving the quadratic program (20) with the same noisy data.
@@
Figure 7(b) shows the final-time electric field $E(x,T)$ for all three operators. The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the central-difference baseline, while the unconstrained LS stencil produces a completely corrupted field. This experiment illustrates a regime where training data are noisy or mildly inconsistent, and the energy constraints act as a form of physics-based regularization: they prevent overfitting to noise and guarantee a discrete Poynting theorem irrespective of the data quality.
(a) Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ under noisy training.
(b) Final-time electric field $E(x,T)$ for noisy training.
Figure 7: Noisy training and ill-posed least squares: comparison of unconstrained LS, central difference, and energy-constrained stencils.
Figure 8: Space-time plot of the electric field $E(x,t)$ under noisy training perturbations.